OpenStudy (anonymous):

Find the limit: lim θ->0 cos θ tanθ/θ

OpenStudy (amistre64):

since cos(0) = 1 this is just a silly little trig game

OpenStudy (amistre64):

unless im reding it wrong; then its a difficult silly little trig game :)

OpenStudy (amistre64):

rewrite as sin(t)/t

OpenStudy (amistre64):

goes to 1

myininaya (myininaya):

\[\cos(\theta) \cdot \frac{\sin(\theta)}{\cos(\theta)}\cdot \frac{1}{\theta} =\frac{\cos(\theta)}{\cos(\theta)} \cdot \frac{\sin(\theta)}{\theta}\]

OpenStudy (amistre64):



OpenStudy (anonymous):

they want you to say "oh i know , this is \[\lim_{\theta \rightarrow 0}\frac{\sin(\theta)}{\theta}\] and i learned that one last week

OpenStudy (amistre64):

as sin(t) and t get close zero; they pretty much converge onto each other so that they go to 1

myininaya (myininaya):

\[\lim_{\theta \rightarrow 0}\frac{\cos(\theta)}{\cos(\theta)} \cdot \frac{\sin(\theta)}{\theta}=\lim_{\theta \rightarrow 0}\frac{\sin(\theta)}{\theta}=1\]

OpenStudy (amistre64):

picture not to scale :)

OpenStudy (anonymous):

\[\color{blue}{\text{amistre }}=\color{red}{\text{picasso}}\]

OpenStudy (amistre64):

there is a "ask a question" box to the top left of the screen; use that to ask questions in, it will get more of an audience that way